# MrT5: Dynamic Token Merging for Efficient Byte-level Language Models

### v2

@@ Appendix H Continued Pre-Training Experiments with Large Models
We present additional continued pre-training experiments using larger model sizes. Specifically, we train one ByT5 model and one MrT5 model, both initialized from the pre-trained 1.23-billion parameter ByT5 Large. While ByT5 Small consists of 12 encoder layers and 4 decoder layers, ByT5 Large has a much deeper architecture with 36 encoder layers and 12 decoder layers. In MrT5, the delete gate is applied at layer $l=3$, and all other data and training configurations remain consistent with those used in our primary continued pre-training experiments, as detailed in Section E.2. We employ a PI-controller to target a deletion ratio of $\delta=0.5$. We evaluate on a smaller batch size of 4 examples ($2^{12}$ tokens) using a test set containing 2500 examples in each of the 15 languages.

### v3

@@ Appendix H Continued Pre-Training Experiments with Large Models
We present additional continued pre-training experiments using larger model sizes. Specifically, we train one ByT5 model and one MrT5 model, both initialized from the pre-trained 1.23-billion parameter ByT5 Large. While ByT5 Small consists of 12 encoder layers and 4 decoder layers, ByT5 Large has a much deeper architecture with 36 encoder layers and 12 decoder layers. ByT5 Large also has $d_{\text{ff}}=3840$, $d_{\text{model}}=1536$, and 16 attention heads in each layer. MrT5 Large’s gating mechanism introduces an additional 3,073 parameters. In MrT5, the delete gate is applied at layer $l=3$, and all other data and training configurations remain consistent with those used in our primary continued pre-training experiments, as detailed in Section E.2. We employ a PI controller to target a deletion ratio of $\delta=0.5$. We evaluate on a smaller batch size of 4 examples ($2^{12}$ tokens) using a test set containing 2500 examples in each of the 15 languages.